# Fast, accurate training and sampling of Restricted Boltzmann Machines

### v1

@@ 7 Conclusions
We have shown that the strategy of initiating the training on a pre-trained low-rank RBM is an extremely effective strategy to obtain high quality models for structured datasets that accurately represent all the modes in the datasets and with significantly higher log-likelihoods. We have also shown that the models obtained in that way are: (i) better generative models than those obtained with standard trainings, both, in the sense that they over-fit less at the same time they are more indistinguishable from the test samples, (ii) they display faster relaxational dynamics.
We have also proposed a new fast sampling method that exploits the progressive learning of features in the training of RBMs to design an efficient trajectory PT strategy that allows accelerating the parallel Gibbs sampling dynamics by many orders of magnitude and overcome the performance of recent efficient sampling methods without adding any extra cost than saving models during the training.
Both strategies for training and sampling are very general, and could be generalized to more complex EBMs. In this sense, the low-rank RBM model could be used as a more efficient pre-initialisation in deeper structures, and the trajectory PT algorithm is suitable to be directly used in any EBM no matter how complex it is.
@@ 8 Code availability
The code and datasets are available at https://github.com/nbereux/fast-RBM.
9 acknowledgments
The authors acknowledge financial support by the Comunidad de Madrid and the Complutense University of Madrid (UCM) through the Atracción de Talento program (Refs. 2019-T1/TIC-13298), and to project PID2021-125506NA-I00 financed by the Ministerio de Economía y Competitividad, Agencia Estatal de Investigación (MICIU/AEI /10.13039/501100011033), and the Fondo Europeo de Desarrollo Regional (FEDER, UE).
@@ A.1 The low-rank RBM and its sampling procedure
The training can therefore be done directly in the subspace of dimension $d$. In Ref. [41], it has been shown that such RBM can be trained by mean of the Restricted Coulomb Machine, where the gradient is actually convex in the parameter’s space. It is then possible to do a mapping from the RCM to the RBM to recover the RBM’s parameters. In brief, the training of the low-dimensional RBM is performed by the RCM, and then the parameters are obtrained via a direct relation between the RCM and the RBM’s parameters. The detail of the definition and of the training of the RCM is detailed in the appendix A.2.
@@ A.2 The Restricted Coulomb Machine
As introduced in [41], it is possible to exactly train a surrogate model for the RBM, called the Restricted Coulomb Machine (RCM), on a low dimensional dataset without explicitly sampling the machine allowing to learn even heavily clustered datasets. We will briefly outline the main steps to train the RCM. A more detailed explanation can be found in Appendix A.2.
@@
$m_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}s_{i}u_{i\alpha},\quad w_{\alpha a}\coloneqq\sum_{i=1}^{N_{v}}w_{ia}u_{i\alpha},\quad\theta_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}\theta_{i}u_{i\alpha}$ (19)
with $\alpha\in\{1,\ldots,d\}$ and $\bm{v}$ the projection matrix of the PCA. The projected distribution of the model is then given by
@@
in order to obtain the RBM from the RCM. The model is then trained through log-likelihood maximization over its parameters. However, this objective is non-convex if all the parameters are trained through gradient ascent. To relax the problem, since we’re in low dimension, we can define a family of hyperplanes $(\bm{n},\bm{z})$ covering the space and let the model only learn the weights of each to the hyperplane. We can then discard the ones with a weight low enough for the approximation (21) to be bad.
@@ Appendix B Sampling via Parallel Tempering using the learning trajectory
Model Dataset # of machines Alg. # of steps acc. factor @ 104 steps
pre-train+PCD MNIST01 6 (+1) 10000 80
JarJar MNIST01 28 10000 50
PCD MNIST01 13 10000 30
pre-train+PCD Human Genome 6 (+1) 10000 1350
PCD Human Genome 13 10000 7100
Table 1: Performance comparison of different models on various datasets for the sampling using PTT versus Gibbs sampling for $10^{4}$ mcmc steps. The acceleration factor is defined as the ratio of the average number of jumps obtained until $10^{4}$ steps between PTT and Gibbs sampling. For pre-train+PCD, the RCM machine has not to be counted among the list of models (hence the +1) because it is very fast to sample from.
@@
A comparison of performances between PTT and standard Gibbs sampling is reported in Tab. 1.
Figure 7: Comparison between PTT and standard Gibbs Sampling for RBMs trained using PCD (A and B) and JarRBM (C and D) on the MNIST01 dataset. A and C show the sampling trajectory of two chains recorded every 10 steps for a total of $10^{4}$ mcmc steps. B and D show the average number of jumps of a population of 100 chains as a function of the sampling time.
Figure 8: Comparison between PTT and standard Gibbs Sampling for RBMs trained using PCD on the Human Genome dataset. The sampling has been performed under the same conditions of fig. 7.
@@ Appendix C Training details
PCD 2000 2000 10 000 0.01 100 185
Jar-RBM 2000 10 000 10 000 0.01 100 185
Pre-train+PCD 2000 2000 10 000 0.01 100 185
@@
PCD 2000 2000 10 000 0.01 100 100
Jar-RBM 2000 10 000 10 000 0.01 100 100
Pre-train+PCD 2000 2000 10 000 0.01 100 100
@@
Appendix D Training of the RBM using the Jarzynski equation
@@
In this section, we describe a procedure similar to the one introduced in [37] for training the RBM by leveraging the Jarzynki equation. In one of its formulations, the Jarzynski equation states that we can relate the ensemble average of an observable $\mathcal{O}$ with the average obtained through many repetitions of an out-of-equilibrium dynamical process. If we consider the training trajectory of an RBM, $p_{0}\rightarrow p_{1}\rightarrow\dots\rightarrow p_{t-1}\rightarrow p_{t}$, we can write
@@
$\langle\mathcal{O}\rangle_{t}=\frac{\left\langle\mathcal{O}e^{-W_{t}}\right\rangle_{\mathrm{traj}}}{\left\langle e^{-W_{t}}\right\rangle_{\mathrm{traj}}},$ (27)
where the average on the lhs is done over the last model $p_{t}$, the averages on the rhs are taken across many different trajectory realizations and $W_{t}$ is a trajectory-dependent importance factor. By all practical means, under the assumption of having quasi-adiabatic parameters updates, namely $p(\Theta_{t-1}\rightarrow\Theta_{t})=p(\Theta_{t}\rightarrow\Theta_{t-1})$, this means that we can assign to each Markov chain of the simulation $\bm{x}^{(r)}$, $r=1,\dots,R,$ an importance weight given by:
$W_{t}^{(r)}=\sum_{\tau=1}^{t}[\mathcal{H}_{\tau}(\bm{x}^{(r)}_{\tau-1})-\mathcal{H}_{\tau-1}(\bm{x}^{(r)}_{\tau-1})]$ (28)
and then compute the gradient of the log-likelihood by means of a weighted average over the chains:
$\langle\mathcal{O}\rangle_{t}\simeq\frac{\sum_{r=1}^{R}\mathcal{O}(\bm{x}^{(r)})~{}e^{-W^{(r)}_{t}}}{\sum_{r=1}^{R}e^{-W^{(r)}_{t}}}.$ (29)
@@
Notice that, since Eq. (27) is an exact result, the importance weights should, in principle, eliminate the bias brought by the non-convergent chains used for approximating the log-likelihood gradient in the classical PCD scheme. However, after many updates of the importance weights, one finds that only a few chains carry almost all the importance mass. In other words, the vast majority of the chains we are simulating are statistically irrelevant, and we expect to get large fluctuations in the estimate of the gradient because of the small effective number of chains contributing to the statistical average. A good observable for monitoring this effect is the Effective Sample Size (ESS), defined as [37]

### v2

@@ 6 Conclusions
In this work, we tackle the challenges of training high-quality equilibrium RBMs on highly structured datasets. We focused on improving the sampling process and accurately evaluating test log-likelihood throughout training. Our findings reveal that for datasets that are tightly clustered in their principal component analysis representation, introducing a pre-training phase can enhance training quality. This phase involves encoding the principal directions of the data into the model using a convex optimization approach. However, this pre-training method, that is specific to RBMs, appears to have limited benefits for less structured datasets or for long trainings, as chains get out of equilibrium, and should be combined in the future with optimized MCMC methods to achieve its maximum capacity.
Furthermore, we developed a new approach that exploits the progressive learning of features during RBM training to enhance evaluation and sampling. This approach enables online and precise computation of log-likelihood and leads to an efficient sampling strategy akin to the standard parallel tempering algorithm. Our results demonstrate that both our proposed methods outperform traditional algorithms without the need for simulating models at unnecessary temperatures or training additional models to facilitate sampling. Furthermore, both the trajectory AIS method for log-likelihood estimation and the trajectory parallel tempering algorithm are highly versatile, making them applicable to more complex energy-based models. Additionally, the low-rank RBM model can serve as an effective pre-initialization step for deeper architectures, and the trajectory PT algorithm is suitable for direct use in any EBM, regardless of complexity.
@@ 7 Code availability
The code and datasets are available at https://github.com/anonymousrbm/iclr-2025
Supplemental Information
@@ A.1 The low-rank RBM and its sampling procedure
The training can therefore be done directly in the subspace of dimension $d$. In Ref. Decelle & Furtlehner (2021a), it has been shown that such RBM can be trained by mean of the Restricted Coulomb Machine, where the gradient is actually convex in the parameter’s space. It is then possible to do a mapping from the RCM to the RBM to recover the RBM’s parameters. In brief, the training of the low-dimensional RBM is performed by the RCM, and then the parameters are obtrained via a direct relation between the RCM and the RBM’s parameters. The detail of the definition and of the training of the RCM is detailed in the appendix A.2.
@@ A.2 The Restricted Coulomb Machine
As introduced in Decelle & Furtlehner (2021a), it is possible to exactly train a surrogate model for the RBM, called the Restricted Coulomb Machine (RCM), on a low dimensional dataset without explicitly sampling the machine allowing to learn even heavily clustered datasets.
@@
$m_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}v_{i}u_{i\alpha},\quad w_{\alpha a}\coloneqq\sum_{i=1}^{N_{v}}w_{ia}u_{i\alpha},\quad\theta_{\alpha}\coloneqq\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}\theta_{i}u_{i\alpha}$ (20)
with $\alpha\in\{0,\ldots,d\}$ and $\bm{u}$ the projection matrix of the PCA.
The projected distribution of the model is then given by
@@
in order to obtain the RBM from the RCM. The model is then trained through log-likelihood maximization over its parameters. However, this objective is non-convex if all the parameters are trained through gradient ascent. To relax the problem, since we’re in low dimension, we can define a family of hyperplanes $(\bm{n},\bm{z})$ covering the space and let the model only learn the weights of each to the hyperplane. We can then discard the ones with a weight low enough for the approximation equation 22 to be bad.
@@
A.3 Learning the bias
To initialize the model with its bias set to the dataset’s mean, we define the 00-th direction of the longitudinal space as the direction obtained by projecting the configuration space onto the normalized mean vector of the dataset.
The RBM will only contribute to this direction through the visible bias:
$H(\bm{m}(v))=-\sum_{\alpha=0}^{N_{v}}\theta_{\alpha}m_{\alpha}-\frac{1}{N_{v}}\sum_{a}\log\cosh\left(\sqrt{N_{v}}\sum_{\alpha=1}^{d}m_{\alpha}w_{\alpha a}+\eta_{a}\right).$ (27)
This implies that this direction is independent of the others within the model. Consequently, we can create an independent mesh for this direction to learn it. The rest of the training is performed using the same procedure as before.
@@ Appendix B Sampling via Parallel Tempering using the learning trajectory
A comparison of performances between PTT and standard Gibbs sampling is reported in Fig.4.
B.1 Controlling equilibration times with PTT
One of the advantages of PT-like approaches is that they allow easy control over thermalization times by studying the random walk in temperatures (or models for the PTT) visited by each run over time, as proposed in Ref. Banos et al. (2010). From now on, we will refer only to the number of models used in the PTT sampling, denoted as $N_{\mathrm{m}}$. In an ergodic sampling process, each run should spend roughly the same amount of time at each of the $N_{\mathrm{m}}$ different model indices, as all models have the same probability.
The mean temperature/model index is known to be $\left\langle n\right\rangle=(N_{\mathrm{m}}-1)/2$ (assuming the model indices run from 00 to $N_{\mathrm{m}}-1$). This simplifies the computation of the time-autocorrelation function of the model indices $n_{i}(t)$ visited by each sample $i$ at time $t$:
$C(t)=\frac{\overline{(n_{i}(t+t_{0})-\left\langle n\right\rangle)(n_{i}(t_{0})-\left\langle n\right\rangle)}}{\overline{(n_{i}(t_{0})-\left\langle n\right\rangle)^{2}}}$
Here, the symbol $\overline{(\cdots)}$ refers to the average over all samples $i$ and initial times $t_{0}$ obtained from our PTT runs. This allows us to compute the exponential autocorrelation time, $\tau_{\mathrm{exp}}$, from a fit to
$C(t)\sim Ae^{-t/\tau_{\mathrm{exp}}}\quad\text{for large }t,$
as well as the integrated autocorrelation time, $\tau_{\mathrm{int}}$, using the self-consistent equation Sokal (1997); Amit & Martin-Mayor (2005):
$\tau_{\mathrm{int}}=\frac{1}{2}+\sum_{t=0}^{6\tau_{\mathrm{int}}}C(t).$
These two times provide different insights: $\tau_{\mathrm{exp}}$ gives insight into the time needed to thermalize the simulations (setting the length of the runs to at least $20\tau_{\mathrm{exp}}$ is generally a safe (and very conservative) choice to ensure thermalization Sokal (1997)), while $\tau_{\mathrm{int}}$ gives the time needed to generate independent samples. To keep track of the fluctuations, we compute a different $C^{m}(t)$ for each $m=1,\ldots,N_{\mathrm{m}}$ system index simulated in parallel in the PTT, where each $C^{m}(t)$ is averaged over 1000 independent realizations (with all runs initialized to the same model index). We then compute the average over all parallel runs $C(t)=\overline{C^{m}(t)}$ and use this curve to extract $\tau_{\mathrm{exp}}$ and $\tau_{\mathrm{int}}$. The errors in $C$ and the times are then extracted from the error of the mean of the values extracted with each of the $C^{m}(t)$ curves.
For example, in Fig. 6 (left), we show the averaged curves $C(t)$ obtained with the PTT sampling of the machines discussed in the main text in the MNIST01 (top) and HGD (bottom) datasets. The extracted times are $\tau_{\mathrm{exp}}=258(29)$ and $\tau_{\mathrm{int}}=9.0(1.5)$ for MNIST01, and $\tau_{\mathrm{exp}}=48(10)$ and $\tau_{\mathrm{int}}=3.8(7)$ for HGD. These numbers guarantee us that the $10^{4}$ steps used to generate the scatterplots in Fig. 2 were long enough to ensure equilibration in this dataset, but also in the others, as shown in Table 1.
Figure 6: (Left) We show the autocorrelation function $C(t)$ obtained as in Eq. equation B.1 when we analyze the random walk in the model indices during PTT sampling with the same RBMs analyzed in the main text (trained with pretraining+PCD) for the MNIST01 (top) and HGD (bottom) datasets. Both curves were obtained with a PTT run of $5\cdot 10^{4}$ MCMC steps, where the models used were automatically selected during training using the PTT moves acceptance criterion of 0.25. (Right) In blue we show the $\tau_{\mathrm{exp}}$ as a function of the target acceptance rate for MNIST01 (top) and HGD (bottom). The error bar results from the error of the mean of the values extracted from the different $C_{m}(t)$ and the shading marks the range between the largest and the smallest value of the $N_{\mathrm{m}}$ curves. In orange, we show the total simulation costs ($\tau_{\mathrm{exp}}$ multiplied by the number of models to be simulated in parallel) as a function of the target acceptance rate. This number of models saved, as a function of the acceptance rate, is shown in green in the inset.
dataset $\tau_{\mathrm{exp}}$ $\tau_{\mathrm{int}}$
MNIST01 258(29) 9.0(15)
HGD 48(10) 3.8(7)
Mickey 2(0) 1(0)
Ising 18(3) 2.4(5)
Table 1: Autocorrelation times associated to the random walk dynamics between models in the PTT sampling obtained with the models trained with pre-training+PCD on different datasets. Models used in the PTT are all chosen with the $a=0.25$ criterion.
With these measures, we can try to optimize the number of models that we use for the PTT and that are automatically selected during training. The criterion here is that the acceptance of PTT exchanges between neighboring models must not go below a target acceptance $a$. In the figures shown so far, $a=0.25$. Let us justify this choice. The total physical time we need to simulate $N_{\mathrm{m}}$ models grows proportionally with $N_{\mathrm{m}}$ (unless one manages to properly parallelize the sampling, which is feasible but we did not do it), while the time needed to flow between very well trained and poorly trained models controlling the mixing time scales as $\sqrt{N_{\mathrm{m}}}/a$, where $a$ is the average acceptance of the exchanges between the models. Since $N_{\mathrm{m}}$ grows with $a$ in practice (as shown in the inset of Fig. 6–right), it is expected that the total time can be optimized for a given value of $a$. In Fig. 6–right we show $\tau_{\mathrm{exp}}$ in blue (and $\tau_{\mathrm{exp}}N_{\mathrm{m}}$ in orange) as a function of $a$ for the data sets MNIST01 (top) and HGD (bottom). These figures show that the choice of $a$ is not particularly crucial as long as it is between 0.2 and 0.5.
@@ Appendix C Training details
PCD 2000 2000 10 000 0.01 100 200
Pre-train+PCD 2000 2000 10 000 0.01 100 200
Ising
PCD 2000 2000 10 000 0.01 100 200
Pre-train+PCD 2000 2000 10 000 0.01 100 200
Mickey
PCD 2000 2000 10 000 0.01 100 200
Pre-train+PCD 2000 2000 10 000 0.01 100 200
@@
MNIST
PCD 2000 2000 10 000 0.01 100 200
Pre-train+PCD 2000 2000 10 000 0.01 100 200
Table 4: Training time of the RBMs
Name Training time (s)
CelebA
PCD $119.2\pm 0.8$
Pre-train+PCD $3134.0\pm 0.6$
HGD
PCD $106.2\pm 0.3$
Pre-train+PCD $266\pm 3$
Ising
PCD $62\pm 1$
Pre-train+PCD $104\pm 2$
@@
PCD $118.3\pm 3$
Pre-train+PCD $246.0\pm 0.7$
MNIST-01
PCD $100.6\pm 0.3$
Pre-train+PCD $258.0\pm 0.9$
MNIST
PCD $105.7\pm 0.2$
Pre-train+PCD $233.9\pm 0.2$
Table 5: RCM Hyperparameters
Dataset name # Dimensions # Hidden nodes Training time (s)
CelebA 4 100 3006
HGD 3 100 150
Ising 3 100 32
Mickey 2 100 118
MNIST-01 3 100 148
MNIST 3 100 119
@@
Appendix D Computation of the likelihood using AIS
To compute the likelihood of the RBMs we used the Annealed Importance Sampling method introduced by Neal (2001) and recently studied in the case of the RBM Krause et al. (2020).
D.1 Offline AIS
We follow the method described in Krause et al. (2020). The AIS estimate is based on creating a sequence of machines, at one end a machine from which we can compute the partition function and at the other end the machine from which we want to compute the partition function. Let’s consider that we have $i=0,\dots,N_{\rm rbm}$ machines described by
$p_{i}(\bm{x})=\frac{e^{-\mathcal{H}_{i}(\bm{x})}}{Z}=\frac{p_{i}^{*}(\bm{x})}{Z},$
where $p_{0}=p_{\rm ref}$ denote the RBM from which we can compute the partition function (typically a RBM where the weight matrix is zero) and $p_{N_{\rm rbm}}=p_{\rm target}$ the machine from which we want to estimate the partition function. For each machine, we denote its set of parameters $\bm{\Theta}_{i}=\{\bm{w}_{i},\bm{\theta}_{i},\bm{\eta}_{i}\}$. Following Krause et al. (2020), defining a transition rule $T_{i}(\bm{x}_{i-1}\to\bm{x}_{i})$ in order to pass from a configuration $\bm{x}_{i-1}$ of the system $i-1$ to $\bm{x}_{i}$ of the system $i$, that we have the following identity
$\frac{Z_{N_{\rm rbm}}}{Z_{0}}=\left\langle\exp\left(-\sum_{i=1}^{N_{\rm rbm}}\mathcal{H}_{i+1}(\bm{x}_{i})-\mathcal{H}_{i}(\bm{x}_{i})\right)\right\rangle_{\rm ForwardPath}.$ (29)
There, if one can compute $Z_{0}$, the estimate of $Z_{N_{\rm rbm}}$ is obtained by averaging the r.h.s. of Eq. 29 over the forward trajectory starting from a configuration $\bm{x}_{0}$ using the operators $T_{i}$. While the method is exact, the variance of the estimation relies on on close are two successive distribution $p_{i}$, $p_{i+1}$.
Temperature AIS:
a commonly used implementation of AIS, is to multiply the Hamiltonian of the system by a temperature, and to the different systems indexed by $i$, by a set of $\beta_{i}$ where $\beta_{0}=0$, $\beta_{i}<\beta_{i+1}$ and $\beta_{N_{\rm rbm}}=1$. Therefore, it is equivalent to choosing $\Theta_{i}=\beta_{i}\{\bm{w},\bm{\theta},\bm{\eta}\}$. Another version of Temperature AIS consist in choosing a slightly different reference distribution $p_{0}$ that take into account the bias of the model. In this setting, the intermediate distribution are taken as
$p_{i}(\bm{x})=\frac{p_{\rm ref}^{(1-\beta_{i})}e^{-\beta_{i}\mathcal{H}(\bm{x})}}{Z_{i}}.$
The main difference is that a prior distribution is used as reference $p_{0}=p_{\rm ref}$ and the temperatures interpolate between this prior and the target distribution. It is quite common to take
$p_{\rm ref}(\bm{x})=\frac{\exp\left(\sum_{i}\theta_{i}v_{i}+\sum_{a}\eta_{a}h_{a}\right)}{\prod_{i}2\cosh(\theta_{i})\prod_{a}2\cosh(\eta_{a})}.$
Trajectory AIS:
the general derivation of AIS allow in principle to choose the set forward probability and/or intermediate system as one sees fit. Since again, a change of temperature in the system can induce discontinuous transition for which, successive distribution will be far away. While, when considering system that are neighbors in the learning trajectory, the system should pass through continuous transition which do not suffer this problem. Hence, we consider a set of RBMs along the learning trajectory. The RBMs are chosen following the prescription that the exchange rate between configuration (see Eq. 5) is about $\sim 0.3$). Then the formula Eq. 29 is used to estimate the partition function.
D.2 Online AIS
The online AIS (Annealed Importance Sampling) is the trajectory AIS evaluated at every gradient update during training. A set of permanent chains is initialized at the beginning of the training process. Every time the parameters are updated, a step is performed on these chains before using them to compute the current importance weights. This approach is computationally efficient relative to the cost of sampling (1 step vs. 100 steps) and allows for a very precise estimation.
Appendix E Training of the RBM using the Jarzynski equation
In this section, we describe the procedure introduced in Carbone et al. (2024) for training energy-based models by leveraging the Jarzynki equation, and we adapt it to the specific case of the RBM. At variance with Carbone et al. (2024), we introduce an additional assumption on the transition probability of the parameters that will lead us to the the reweighting implemented in population annealing methods Weigel et al. (2021).
@@ Appendix E Training of the RBM using the Jarzynski equation
In one of its formulations, the Jarzynski equation states that we can relate the ensemble average of an observable $\mathcal{O}$ with the average obtained through many repetitions of an out-of-equilibrium dynamical process. Let us consider a path $\boldsymbol{y}=\{(\boldsymbol{x}^{0},\boldsymbol{\theta}^{0}),(\boldsymbol{x}^{1},\boldsymbol{\theta}^{1}),\dots,(\boldsymbol{x}^{t},\boldsymbol{\theta}^{t})\}$, where $\bm{x}^{i}=(\bm{v}^{i},\bm{h}^{t})$ refers to the model variables and $\bm{\theta}^{i}=(w^{i},\eta^{i},\theta^{i})$, the model parameters, both at time $i$.
If we consider the training trajectory of an RBM, $p_{0}\rightarrow p_{1}\rightarrow\dots\rightarrow p_{t-1}\rightarrow p_{t}$, we can write the average of an observable $\mathcal{O}$ over the last model of the trajectory $p_{t}$ as
@@
$\langle\mathcal{O}\rangle_{t}=\frac{\left\langle\mathcal{O}~{}e^{-W^{t}}\right\rangle_{\mathrm{traj}}}{\left\langle e^{-W^{t}}\right\rangle_{\mathrm{traj}}}=\frac{\sum_{r=1}^{R}\mathcal{O}(\boldsymbol{y}_{r})~{}e^{-W_{r}^{t}}}{\sum_{r=1}^{R}e^{-W_{r}^{t}}},$ (30)
where $W^{t}$ is a trajectory-dependent importance factor and the averages on the rhs are taken across many different trajectory realizations. In particular, $W_{t}$ is given by the recursive formula:
$W^{t}(\bm{y})$ $=$ $\mathcal{H}^{t}(\bm{x}^{t})-\mathcal{H}^{0}(\bm{x}^{0})-\sum_{i=1}^{t}\log\tilde{T}^{i}(\bm{y}^{i}\rightarrow\bm{y}^{i-1})+\sum_{i=1}^{t}\log T^{i}(\bm{y}^{i-1}\rightarrow\bm{y}^{i})$ (31)
$=$ $W^{t-1}(\bm{y})+\mathcal{H}^{t}(\bm{x}^{t})-\mathcal{H}^{t-1}(\bm{x}^{t-1})-\log\frac{\tilde{T}^{t}(\bm{y}^{t}\rightarrow\bm{y}^{t-1})}{T^{t}(\bm{y}^{t-1}\rightarrow\bm{y}^{t})}$ (32)
Where $T$ and $\tilde{T}$ are, respectively, the forward and backward transition probabilities.
The transition matrix between two states $\bm{y}^{t-1}\rightarrow\bm{y}^{t}$ is factorized into two different subsequent movements:
  1. 1.
    Update the variables $\bm{x}^{t-1}\to\bm{x}^{t}$ on the fixed model $\bm{\theta}^{t-1}$,
  2. 2.
    Update the parameters $\bm{\theta}^{t-1}\to\bm{\theta}^{t}$ on fixed variables $\bm{x}^{t}$.
In other words,
$T(\bm{y}^{t-1}\to\bm{y}^{t})=w_{\bm{\theta}^{t-1}}(\bm{x}^{t-1}\to\bm{x}^{t})t_{\bm{x}^{t}}(\bm{\theta}^{t-1}\to\bm{\theta}^{t}),$ (33)
and
$\tilde{T}(\bm{y}^{t}\to\bm{y}^{t-1})=t_{\bm{x}^{t}}(\bm{\theta}^{t}\to\bm{\theta}^{t-1})w_{\bm{\theta}^{t-1}}(\bm{x}^{t}\to\bm{x}^{t-1}),$ (34)
so that
$\log\frac{\tilde{T}^{t}(\bm{y}^{t}\rightarrow\bm{y}^{t-1})}{T^{t}(\bm{y}^{t-1}\rightarrow\bm{y}^{t})}=\log\frac{w_{\bm{\theta}^{t-1}}(\bm{x}^{t}\to\bm{x}^{t-1})}{w_{\bm{\theta}^{t-1}}(\bm{x}^{t-1}\to\bm{x}^{t})}+\log\frac{t_{\bm{x}^{t}}(\bm{\theta}^{t}\to\bm{\theta}^{t-1})}{t_{\bm{x}^{t}}(\bm{\theta}^{t-1}\to\bm{\theta}^{t})}.$ (35)
We know the ratio between the transition probabilities between $\bm{x}^{t-1}\to\bm{x}^{t}$, because evolution of the variables during the simulation is performed using MCMC simulations that satisfy detailed balance. This means that:
$\log\frac{w_{\bm{\theta}^{t-1}}(\bm{x}^{t}\to\bm{x}^{t-1})}{w_{\bm{\theta}^{t-1}}(\bm{x}^{t-1}\to\bm{x}^{t})}=\mathcal{H}^{t-1}(\bm{x}^{t})-\mathcal{H}^{t-1}(\bm{x}^{t-1}),$ (36)
and we are left with
$W^{t}(\bm{y})$ $=$ $W^{t-1}(\bm{x})+\mathcal{H}^{t}(\bm{x}^{t})-\mathcal{H}^{t-1}(\bm{x}^{t})-\log\frac{t_{\bm{x}^{t}}(\bm{\theta}^{t}\to\bm{\theta}^{t-1})}{t_{\bm{x}^{t}}(\bm{\theta}^{t-1}\to\bm{\theta}^{t})}.$ (37)
The standard formula that we use in the paper assumes that the changes in $\bm{\theta}$ are so small that we can consider the transition probabilities to be symmetrical and the last term in the rhs cancels out. In that case, we recover the known expressions of the population annealing.
@@
Notice that, since Eq. equation 30 is an exact result, the importance weights should, in principle, eliminate the bias brought by the non-convergent chains used for approximating the log-likelihood gradient in the classical PCD scheme. However, after many updates of the importance weights, one finds that only a few chains carry almost all the importance mass. In other words, the vast majority of the chains we are simulating are statistically irrelevant, and we expect to get large fluctuations in the estimate of the gradient because of the small effective number of chains contributing to the statistical average. A good observable for monitoring this effect is the Effective Sample Size (ESS), defined as Carbone et al. (2024)